# Copula‐Stein Discrepancy: A Generator‐Based Stein Operator for Archimedean Dependence

### v1

@@ 7.3 Examples of Non-Archimedean Copulas
Example 7.1 (Elliptical Copulas: Gaussian case)
@@
$\log c_{\Sigma}(u)=-\tfrac{1}{2}\,z(u)^{\top}\!\big(\Sigma^{-1}-I\big)\,z(u)-\tfrac{1}{2}\log\det\Sigma.$
@@
$s_{j}(u)=\frac{\partial\log c_{\Sigma}(u)}{\partial u_{j}}=-\,\frac{\big[(\Sigma^{-1}-I)\,z(u)\big]_{j}}{\varphi_{\mathrm{N}}\!\big(z_{j}(u)\big)},\qquad z_{j}(u)=\Phi^{-1}(u_{j}),$
where $\varphi_{\mathrm{N}}(x)=(2\pi)^{-1/2}\exp(-x^{2}/2)$ is the standard normal PDF (to avoid collision with the random feature map $\phi$ in Sec. 8) and $\cdot]_{j}$ extracts the $j$-th component.
This score satisfies the growth conditions in Assumption 1.
Example 7.2 (Vine Copulas)
@@
Example 7.3 (Mixture Copulas)
@@
where $w_{k}>0$, $\sum_{k}w_{k}=1$, and each $C_{k}$ satisfies Assumption 1.
@@
If each component copula $C_{k}$ satisfies the conditions in Assumption 1, then the mixture also satisfies them.
7.4 Computational Considerations
@@
Proposition 7.1 (Computational Complexity for General Copulas)
@@
Proof. The kernel evaluation formula is identical to the Archimedean case (Theorem 8), so the complexity analysis follows immediately from Proposition 5.4.
@@ 7.5 Summary of the Extension
8.  Computational Complexity and Fast Algorithms
@@
8.1 Exact Complexity Analysis
@@
Theorem 11 (Detailed Complexity Analysis)
@@
Proof. From the general kernel formula (Theorem 8):
@@
Overall Complexity: $O(d)+O(d)+O(d)=O(d)$ operations per kernel evaluation.
8.2 Efficient Algorithms
We now present explicit algorithms for computing CSD with optimal complexity.
Below we have Algorithm 1 for efficient Copula-Stein kernel evaluation.
@@ 8.2 Efficient Algorithms
[…]
@@
1:Sample $\{U_{1},\ldots,U_{n}\}\subset[0,1]^{d}$; target copula parameters
[…]
@@
6:   $k\leftarrow\textsc{CSDKernel}(U_{i},U_{j})$ $\triangleright$ Algorithm 1
@@
11:return $S$
Proposition 8.1 (Algorithm Correctness and Complexity)
@@
Proof. Correctness: Algorithm 1 implements the exact formula from Theorem 8. The key computational insight is in steps 8-14, where we compute the gradient of the product kernel using the identity:
@@
Total: $O(d)$ per kernel evaluation, $O(n^{2}d)$ for full CSD.
@@
8.3 Random Feature Approximations for Scalable CSD
@@
Setup. Let $\phi:[0,1]^{d}\to\mathbf{R}^{m}$ be a random feature map for the base kernel $k$, satisfying
@@
and denote by $\nabla_{u}\phi(u)\in\mathbf{R}^{m\times d}$ its Jacobian. Recall the witness‐function in the RKHS,
@@
Definition 8.1 (Random‐Feature Witness)
@@
$\nabla_{u}\phi(u)+\phi(u)\,s(u)^{\top}\;\in\;\mathbf{R}^{m\times d}$ (79)
@@
and $\mathrm{vec}(\cdot)$ stacks its columns into a vector in $\mathbf{R}^{dm}$.
Assumption 3 (RF smoothness)
@@
$\mathbb{E}_{\phi}[\partial_{u_{j}}\phi(u)^{\top}\phi(v)]=\partial_{u_{j}}k(u,v),\quad\mathbb{E}_{\phi}[\partial_{u_{j}}\phi(u)^{\top}\partial_{v_{j}}\phi(v)]=\partial_{u_{j}}\partial_{v_{j}}k(u,v).$
@@
Theorem 12 (Unbiased RF–CSD)
Under Assumption 3, define $\tilde{\Xi}_{C_{0}}(u)=\mathrm{vec}\!\left(\nabla_{u}\phi(u)+\phi(u)\,s_{0}(u)^{\top}\right)\in\mathbf{R}^{dm}$. Then for any sample $U_{1:n}$,
$\mathbb{E}_{\phi}\!\left[\Big\|\frac{1}{n}\sum_{i=1}^{n}\tilde{\Xi}_{C_{0}}(U_{i})\Big\|_{2}^{2}\,\bigg|\,U_{1:n}\right]=\frac{1}{n^{2}}\sum_{i,j=1}^{n}k_{C_{0}}(U_{i},U_{j}).$
Proof. Since
$\widehat{D}_{n,\text{RF}}^{2}=\Bigl\|\frac{1}{n}\sum_{i}\tilde{\Xi}_{C}(U_{i})\Bigr\|_{2}^{2}=\frac{1}{n^{2}}\sum_{i,j}\tilde{\Xi}_{C}(U_{i})^{\top}\tilde{\Xi}_{C}(U_{j}),$ (81)
@@
$\mathbb{E}_{\phi}\bigl[\widehat{D}_{n,\text{RF}}^{2}\mid U_{1:n}\bigr]=\frac{1}{n^{2}}\sum_{i,j}\mathbb{E}_{\phi}\bigl[\tilde{\Xi}_{C}(U_{i})^{\top}\tilde{\Xi}_{C}(U_{j})\bigr].$ (82)
@@
which is exactly the Copula‐Stein kernel $k_{C_{0}}(u,v)$ by Proposition 3.7. Substituting back into the conditional expectation completes the proof. Finally, computing each $\tilde{\Xi}_{C}(U_{i})$ requires $O(md)$ to form $\phi(U_{i})$ and $\nabla_{u}\phi(U_{i})$ plus an $O(md)$ combine with $s(U_{i})$, totaling $O(nmd)$.
@@
[…]
@@
8:  $J\leftarrow-\sqrt{2/m}\,\mathrm{diag}(\sin z)\,W\in\mathbb{R}^{m\times d}$ $\triangleright$ $J=\nabla_{u}\phi(u)\big|_{u=U_{i}}$
9:  $G\leftarrow J+\phi S^{\top}\in\mathbb{R}^{m\times d}$ $\triangleright$ Stein feature matrix
@@
13:return $\widehat{\mathrm{CSD}}_{\mathrm{RF}}$
@@ 8.4 Parallelization and GPU Implementation
Proposition 8.2 (Parallel Complexity)
@@
Proof. The kernel matrix computation is embarrassingly parallel since each $k_{C}(U_{i},U_{j})$ is independent. Modern GPUs can achieve $P=O(10^{3})$ parallel threads, making large-scale computation feasible.
For the random feature approach, the bottleneck becomes feature computation ($O(nmd)$ total work), which also parallelizes perfectly across samples.
@@
1:Sample $U_{1:n}$ in GPU global memory
2:Full kernel matrix $K\in\mathbb{R}^{n\times n}$ in GPU global memory
[…]
@@
8:  $u\leftarrow U_{i}$ $\triangleright$ Load to thread-local registers
@@
10:  $K[i,j]\leftarrow\textsc{CSDKernel}(u,v)$ $\triangleright$ Uses Alg. 1
@@
1:Kernel matrix $K\in\mathbb{R}^{n\times n}$ in device (GPU) global memory; launch with $B$ blocks, $T$ threads per block; workspace $P\in\mathbb{R}^{B}$ in device memory
[…]
@@
9:  $i\leftarrow p\div n$; $j\leftarrow p\bmod n$
@@
22:  $P[\text{blockIdx}.x]\leftarrow s[0]$ $\triangleright$ Block partial sum to global memory
@@
27:return $S/n^{2}$ $\triangleright$ This is $\widehat{\mathrm{CSD}}_{n}^{2}$
8.5 Memory Optimization
@@
Proposition 8.3 (Memory-Efficient CSD)
@@
Below, we provide the Algorithm 6 for memory-efficient streaming CSD.
@@
1:Sample $U_{1:n}\subset[0,1]^{d}$, block size $B$, kernel routine CSDKernel(,)($\cdot$$\cdot$)
@@
[…]
6:for $p\leftarrow 0$ to M1M-1 do
7:  $i_{\text{lo}}\leftarrow pB$; $i_{\text{hi}}\leftarrow\min\{(p{+}1)B,\,n\}-1$
8:  for $q\leftarrow 0$ to M1M-1 do $\triangleright$ use $q\leftarrow p$ to M1M{-}1 if exploiting symmetry
9:   $j_{\text{lo}}\leftarrow qB$; $j_{\text{hi}}\leftarrow\min\{(q{+}1)B,\,n\}-1$
[…]
11:   $S\leftarrow 0$ $\triangleright$ block partial sum
12:   for $i\leftarrow i_{\text{lo}}$ to $i_{\text{hi}}$ do
13:     for $j\leftarrow j_{\text{lo}}$ to $j_{\text{hi}}$ do
14:      $x\leftarrow\textsc{CSDKernel}(U_{i},U_{j})$
15:      $y\leftarrow x-c$;  $t\leftarrow S+y$;  $c\leftarrow(t-S)-y$;  $S\leftarrow t$ $\triangleright$ Kahan add
16:     end for
17:   end for
18:   if not using symmetry then
19:     $T\leftarrow T+S$
20:   else
21:     if $p=q$ then
22:      $T\leftarrow T+S$ $\triangleright$ diagonal block counts once
23:     else
24:      $T\leftarrow T+2S$ $\triangleright$ off-diagonal mirrored
25:     end if
26:   end if
27:  end for
28:end for
29:return $T/n^{2}$
8.6 Empirical Performance Analysis
We provide theoretical predictions for practical performance based on our complexity analysis.
Proposition 8.4 (Performance Scaling)
For typical implementations on modern hardware:
  1. 1.
    CPU (single-threaded): $T(n,d)\approx c_{1}n^{2}d$ seconds for $c_{1}\approx 10^{-8}$
  2. 2.
    CPU (multi-threaded): $T(n,d)\approx c_{1}n^{2}d/P$ for $P$ cores
  3. 3.
    GPU: $T(n,d)\approx c_{2}n^{2}d/P_{GPU}+c_{3}nd$ for $c_{2}\approx 10^{-10}$, $c_{3}\approx 10^{-6}$
  4. 4.
    Random Features: $T(n,d,m)\approx c_{4}nmd$ for $c_{4}\approx 10^{-7}$
These constants are derived from micro-benchmarks of the basic operations (floating-point arithmetic, memory access, function evaluation) on representative hardware.
8.7 Summary of Computational Viability
Our computational analysis confirms that CSD is a practical tool for modern data science. We have shown that the exact kernel evaluation scales linearly with dimension, which is optimal. For the challenge of large sample sizes, we developed a theoretically-sound random feature approximation that achieves a nearly linear $O(nmd)$ complexity. Finally, we outlined how both the exact and approximate methods can be dramatically accelerated using parallel hardware (Nickolls et al., 2008) and optimized for memory-constrained environments. This combination of theoretical guarantees and computational efficiency makes CSD a powerful and scalable method for real-world applications.
9.  Conclusion and Future Work
9.1 Conclusion
In this work, we introduced the Copula-Stein Discrepancy (CSD), a novel framework that fundamentally shifts the paradigm of goodness-of-fit testing from joint-density-based methods to dependence-structure-aware approaches. By constructing a Stein operator directly on copula densities, CSD addresses a critical blind spot in existing kernel-based methods: their insensitivity to higher-order dependence structures that are crucial in many scientific and practical domains.
Theoretical Contributions. Our theoretical analysis provides the first comprehensive treatment of Stein methods in the copula domain. We established that CSD: (i) metrizes weak convergence on copula distributions, ensuring it detects any mismatch in dependence structure; (ii) achieves the minimax optimal convergence rate of $O_{P}(n^{-1/2})$; and (iii) is provably sensitive to tail dependence differences, with explicit lower bounds on the discrepancy.
Methodological Advances. For Archimedean copulas, we derived closed-form expressions for the CSD kernel. Crucially, we proved that our framework extends far beyond this special case to any copula satisfying mild regularity conditions, encompassing elliptical, vine, and mixture copulas. This generality positions CSD as a unified tool for dependence modeling.
Computational Innovation. We addressed the practical challenges of kernel methods through multiple algorithmic contributions, including exact computation with linear scaling in dimension and random feature approximations that reduce sample complexity from $O(n^{2})$ to nearly linear time. These advances make CSD viable for modern large-scale applications.
CSD represents more than a new statistical test—it embodies a paradigm shift toward structure-aware inference. This paper has established the complete theoretical foundation and the practical, scalable algorithms necessary for dependence-centric model diagnostics in the era of complex, high-dimensional data.
9.2 Future Work
The theoretical framework established in this paper opens several high-impact research directions that span methodology, applications, and long-term vision:
9.2.1 Immediate Extensions
Two-Sample Testing. Developing a CSD-based two-sample test for comparing dependence structures would address a fundamental need in exploratory data analysis.
Parameter Estimation. Using CSD as an objective function for copula parameter estimation presents a compelling, robust alternative to maximum likelihood, particularly for tail dependence parameters.
Model Selection. Extending CSD to select between competing copula families offers a principled approach to structural model selection based on discrepancy minimization.
9.2.2 Methodological and Theoretical Innovations
Deep Generative Models. Incorporating CSD as a regularization term in VAEs or GANs could improve their ability to capture complex dependencies.
High-Dimensional Theory. Extending our analysis to the regime where dimension $d$ grows with sample size $n$ is an important next step.
Streaming Algorithms. Developing online versions of CSD for streaming data would enable real-time dependence monitoring in applications like algorithmic trading.
9.2.3 Empirical Validation and Applications
While this work establishes the theoretical foundations, systematic empirical validation represents a crucial future step. We envision several key directions:
Benchmark Studies. Comprehensive simulation studies comparing CSD against established copula goodness-of-fit tests would be needed to quantify its practical advantages in finite samples, with a focus on power to detect tail dependence misspecification.
Real-Data Applications. The framework can be validated in domains where tail dependence is critical, such as financial risk modeling, hydrology, and neuroscience, to demonstrate its ability to reveal insights missed by traditional methods.
Software Implementation. The development of open-source implementations in R and Python will be essential for facilitating adoption and enabling reproducible research.
9.2.4 Long-Term Vision
Automated Dependence Discovery. In the long term, CSD could form the core of algorithms that automatically discover and characterize dependence structures in complex datasets.
Integration with Causal Inference. Exploring connections between CSD and causal discovery could lead to new approaches for identifying causal relationships based on changes in dependence structure.
Quantum Extensions. As quantum machine learning develops, extending CSD to quantum probability distributions could provide new tools for quantum state verification and dependence modeling.

### v2

@@ B.3 Examples of Non-Archimedean Copulas
Example B.5 (Elliptical Copulas: Gaussian case).
@@
$\log c_{\Sigma}(u)=-\tfrac{1}{2}\,z(u)^{\top}\!\big(\Sigma^{-1}-I\big)\,z(u)-\tfrac{1}{2}\log\det\Sigma.$ (104)
@@
$s_{j}(u)=\frac{\partial\log c_{\Sigma}(u)}{\partial u_{j}}=-\,\frac{\big[(\Sigma^{-1}-I)\,z(u)\big]_{j}}{\varphi_{\mathrm{N}}\!\big(z_{j}(u)\big)},\qquad z_{j}(u)=\Phi^{-1}(u_{j}),$ (105)
where $\varphi_{\mathrm{N}}(x)=(2\pi)^{-1/2}\exp(-x^{2}/2)$ is the standard normal PDF (to avoid collision with the random feature map $\phi$ in Sec. C) and $\cdot]_{j}$ extracts the $j$-th component.
This score satisfies the growth conditions in Assumption 4.2.
Example B.6 (Vine Copulas).
@@
Example B.7 (Mixture Copulas).
@@
where $w_{k}>0$, $\sum_{k}w_{k}=1$, and each $C_{k}$ satisfies Assumption 4.2.
@@
If each component copula $C_{k}$ satisfies the conditions in Assumption 4.2, then the mixture also satisfies them.
B.4 Computational Considerations
@@
Proposition B.8 (Computational Complexity for General Copulas).
@@
Proof.
The kernel evaluation formula is identical to the Archimedean case (Theorem B.2), so the complexity analysis follows immediately from follows immediately from Theorem C.1. ∎
@@ B.5 Summary of the Extension
Appendix C Computational Complexity and Algorithms
@@
C.1 Exact Complexity Analysis
@@
Theorem C.1 (Detailed Complexity Analysis).
@@
Proof.
From the general kernel formula (Theorem B.2):
@@
Overall Complexity: $O(d)+O(d)+O(d)=O(d)$ operations per kernel evaluation. ∎
Below, we give Algorithm 1 for efficient Copula-Stein kernel evaluation.
@@
1:Input: $u,v\in[0,1]^{d}$; Archimedean generator $\varphi$ with $\varphi^{\prime},\varphi^{\prime\prime},\varphi^{(d)},\varphi^{(d+1)}$ and inverse $\varphi^{-1}$; base 1D kernel $k_{1}$ with partials $\partial_{1}k_{1},\partial_{2}k_{1},\partial_{12}k_{1}$
2:Output: $k_{C}(u,v)$
3:Product base kernel and needed derivatives
@@
1:Input: Sample $\{U_{1},\ldots,U_{n}\}\subset[0,1]^{d}$; target copula parameters
2:Output: $\widehat{\mathrm{CSD}}_{n}^{2}$
@@
6:   $k\leftarrow\textsc{CSDKernel}(U_{i},U_{j})$ {Algorithm 1}
@@
11:Return: $S$
Proposition C.2 (Algorithm Correctness and Complexity).
@@
Proof.
Correctness: Algorithm 1 implements the exact formula from Theorem B.2. The key computational insight is in steps 8-14, where we compute the gradient of the product kernel using the identity:
@@
Total: $O(d)$ per kernel evaluation, $O(n^{2}d)$ for full CSD. ∎
@@
C.2 Random Feature Approximations for Scalable CSD
@@
Setup. Let $\phi:[0,1]^{d}\to\mathbb{R}^{m}$ be a random feature map for the base kernel $k$, satisfying
@@
and denote by $\nabla_{u}\phi(u)\in\mathbb{R}^{m\times d}$ its Jacobian. Recall the witness‐function in the RKHS,
@@
Definition C.3 (Random‐Feature Witness).
@@
$\nabla_{u}\phi(u)+\phi(u)\,s(u)^{\top}\;\in\;\mathbb{R}^{m\times d}$ (123)
@@
and $\mathrm{vec}(\cdot)$ stacks its columns into a vector in $\mathbb{R}^{dm}$.
Assumption C.4 (RF smoothness).
@@
$\mathbb{E}_{\phi}[\partial_{u_{j}}\phi(u)^{\top}\phi(v)]=\partial_{u_{j}}k(u,v),\quad\mathbb{E}_{\phi}[\partial_{u_{j}}\phi(u)^{\top}\partial_{v_{j}}\phi(v)]=\partial_{u_{j}}\partial_{v_{j}}k(u,v).$ (125)
@@
Theorem C.5 (Unbiased RF–CSD).
Under Assumption C.4, define $\tilde{\Xi}_{C_{0}}(u)=\mathrm{vec}\!\left(\nabla_{u}\phi(u)+\phi(u)\,s_{0}(u)^{\top}\right)\in\mathbb{R}^{dm}$. Then for any sample $U_{1:n}$,
$\mathbb{E}_{\phi}\!\left[\Big\|\frac{1}{n}\sum_{i=1}^{n}\tilde{\Xi}_{C_{0}}(U_{i})\Big\|_{2}^{2}\,\bigg|\,U_{1:n}\right]=\frac{1}{n^{2}}\sum_{i,j=1}^{n}k_{C_{0}}(U_{i},U_{j}).$ (126)
Proof.
Since
$\widehat{D}_{n,\mathrm{RF}}^{2}=\Bigl\|\frac{1}{n}\sum_{i}\tilde{\Xi}_{C}(U_{i})\Bigr\|_{2}^{2}=\frac{1}{n^{2}}\sum_{i,j}\tilde{\Xi}_{C}(U_{i})^{\top}\tilde{\Xi}_{C}(U_{j}),$ (127)
@@
$\mathbb{E}_{\phi}\bigl[\widehat{D}_{n,\mathrm{RF}}^{2}\mid U_{1:n}\bigr]=\frac{1}{n^{2}}\sum_{i,j}\mathbb{E}_{\phi}\bigl[\tilde{\Xi}_{C}(U_{i})^{\top}\tilde{\Xi}_{C}(U_{j})\bigr].$ (128)
@@
which is exactly the Copula‐Stein kernel $k_{C_{0}}(u,v)$ by Proposition 3.7. Substituting back into the conditional expectation completes the proof. Finally, computing each $\tilde{\Xi}_{C}(U_{i})$ requires $O(md)$ to form $\phi(U_{i})$ and $\nabla_{u}\phi(U_{i})$ plus an $O(md)$ combine with $s(U_{i})$, totaling $O(nmd)$. ∎
@@
1:Input: $U_{1:n}\subset[0,1]^{d}$; RF dimension $m$; bandwidth $\sigma$; score oracle $s(u)\in\mathbb{R}^{d}$
2:Output: $\widehat{\mathrm{CSD}}_{\mathrm{RF}}$
@@
8:  $J\leftarrow-\sqrt{2/m}\,\mathrm{diag}(\sin z)\,W\in\mathbb{R}^{m\times d}$ $J=\nabla_{u}\phi(u)\big|_{u=U_{i}}$}
9:  $G\leftarrow J+\phi S^{\top}\in\mathbb{R}^{m\times d}$ {Stein feature matrix}
@@
13:Return: $\widehat{\mathrm{CSD}}_{\mathrm{RF}}$
@@ C.3 Parallelization and GPU Implementation
Proposition C.6 (Parallel Complexity).
@@
Proof.
The kernel matrix computation is embarrassingly parallel since each $k_{C}(U_{i},U_{j})$ is independent. Modern GPUs can achieve $P=O(10^{3})$ parallel threads, making large-scale computation feasible.
For the random feature approach, the bottleneck becomes feature computation ($O(nmd)$ total work), which also parallelizes perfectly across samples. ∎
@@
1:Input: Sample $U_{1:n}$ in GPU global memory
2:Output: Full kernel matrix $K\in\mathbb{R}^{n\times n}$ in GPU global memory
3: {Kernel launch uses $n\times n$ threads arranged in a 1D or 2D grid}
@@
8:  $u\leftarrow U_{i}$ {Load to thread-local registers}
@@
10:  $K[i,j]\leftarrow\textsc{CSDKernel}(u,v)$ {Uses Alg. 1}
@@
1:Input: Kernel matrix $K\in\mathbb{R}^{n\times n}$ in device (GPU) global memory; launch with $B$ blocks, $T$ threads per block; workspace $P\in\mathbb{R}^{B}$ in device memory
2:Output: $\widehat{\mathrm{CSD}}_{n}^{2}=\big(\sum_{i,j}K[i,j]\big)/n^{2}$ on host (CPU)
3: {Device kernel (block-wise reduction)}
@@
9:  $i\leftarrow\left\lfloor p/n\right\rfloor$; $j\leftarrow p\bmod n$
@@
22:  $P[\text{blockIdx}.x]\leftarrow s[0]$ {Block partial sum to global memory}
@@
27:Return: $S/n^{2}$ {This is $\widehat{\mathrm{CSD}}_{n}^{2}$}
C.4 Memory Optimization
@@
Proposition C.7 (Memory-Efficient CSD).
@@
The Algorithm 6 for memory-efficient streaming CSD is given below.
@@
1:Input: Sample $U_{1:n}\subset[0,1]^{d}$, block size $B$, kernel routine CSDKernel(,)($\cdot$$\cdot$)
2:Output: $\widehat{\mathrm{CSD}}_{n}^{2}=T/n^{2}$ where $K[i,j]=\textsc{CSDKernel}(U_{i},U_{j})$
@@
6: {Option: to exploit symmetry, set $q\leftarrow p$ to M1M{-}1 and accumulate accordingly}
7: {Assume a boolean flag $\textit{useSymmetry}\in\{\texttt{true},\texttt{false}\}$ is set externally}
8:for $p\leftarrow 0$ to M1M-1 do
9:  $i_{\text{lo}}\leftarrow pB$; $i_{\text{hi}}\leftarrow\min\{(p{+}1)B,\,n\}-1$
10:  for $q\leftarrow 0$ to M1M-1 do
11:   $j_{\text{lo}}\leftarrow qB$; $j_{\text{hi}}\leftarrow\min\{(q{+}1)B,\,n\}-1$
12:   {Optionally prefetch $U_{i_{\text{lo}}:i_{\text{hi}}}$ and $U_{j_{\text{lo}}:j_{\text{hi}}}$ into a small buffer}
13:   $S\leftarrow 0$ {block partial sum}
14:   for $i\leftarrow i_{\text{lo}}$ to $i_{\text{hi}}$ do
15:    for $j\leftarrow j_{\text{lo}}$ to $j_{\text{hi}}$ do
16:     $x\leftarrow\textsc{CSDKernel}(U_{i},U_{j})$
17:     $y\leftarrow x-c$;  $t\leftarrow S+y$;  $c\leftarrow(t-S)-y$;  $S\leftarrow t$ {Kahan add}
18:    end for
19:   end for
20:   if not useSymmetry then
21:    $T\leftarrow T+S$
22:   else
23:    if $p=q$ then
24:     $T\leftarrow T+S$ {diagonal block counts once}
25:    else
26:     $T\leftarrow T+2S$ {off-diagonal mirrored}
27:    end if
28:   end if
29:  end for
30:end for
31:Return: $T/n^{2}$
C.5 Computational Scaling Analysis
We summarize the computational costs implied by the CSD V-statistic and by the proposed scalable approximations. The exact computation is quadratic in sample size due to pairwise kernel evaluations, while random-feature approximations reduce the dependence on $n$.
Proposition C.8 (Scaling of computational cost).
Assume the evaluation of the Stein kernel $k_{C_{0}}(u,v)$ costs $O(d)$ per pair $(u,v)\in(0,1)^{d}\times(0,1)^{d}$. Then the V-statistic estimator
$\widehat{\mathrm{CSD}}_{n}^{2}:=\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}k_{C_{0}}(U_{i},U_{j})$ (133)
can be computed in $O(n^{2}d)$ time and $O(1)$ additional memory beyond storing the data.
Moreover, for a random-feature approximation based on $m$ features, the corresponding estimator can be computed in $O(nmd)$ time (and $O(m)$ memory), yielding nearly linear scaling in $n$ when $m\ll n$.
C.6 Summary of computational viability
The above scaling highlights a standard trade-off. Exact CSD computation is feasible for moderate $n$ and benefits from vectorization and parallelism, while the random-feature approximation provides a practical route for large sample sizes by replacing the quadratic dependence on $n$ with a dependence on the chosen feature budget $m$. In practice, both exact and approximate computations can be accelerated via standard parallel hardware and implementation-level optimizations (Nickolls et al., 2008).